How far upstream can a canoer travel at 8 mi/hr and still return downstream at 12mi/hr in a total of 4 hours?
Let t = the time to travel upstream.
4-t = the time to travel downstream.
Ideal for equation:
Distance traveled upstream = Distance traveled downstream

Answer by mananth(16946)

let the distance the canoe can travel bex miles
..
speed downstream = 12mph
speed upstream = 8 mph.
...
time = distance /speed
...
Time upstream + time downstream = 4 hours
x/8 + x/12 =4
LCM = 24
(3x+2x)/24 = 4
5x/24 =4
5x=96
x= 19.2 miles
....
C H E C K
19.2/8 +19.2/12
2.4 + 1.6 = 4 hours
...
m.ananth@hotmail.ca